1. $800 invested at an annually compounded interest rate of 6 percent will be worth how much at the end of 10 years?

a) $1,280.

b) $1,433.

c) $1,417.

d) $1,369.

2. What is the present value of $500 to be received eight years from now if the interest rate is 5 percent?

a) $300.

b) $338.42.

c) $700.

d) $738.72

3. Kira has $2,000 to invest today that she wants to grow to $3,000 in five years. What annually compounded rate of interest would she have to receive in order to reach her goal?

a) Roughly 4.6 percent.

b) Roughly 6.5 percent.

c) Roughly 8.4 percent.

d) Roughly 9.3 percent.
Answer #1

1. Ans: b) $1433

Explanation:

F = P ( F / P , i ,n )

= 800 ( F/P , 6% , 10 )

= 800 ( 1.791 )

= $1432.8

2. Ans: b ) $338.42

Explanation:

P = F ( P / F , i ,n )

= 500 ( P /F , 5% , 8 )

= 500 * 0.6768

= $338.40

3. Ans: c ) Roughly 8.4 percent

Explanation:

F = P( 1+ i)n

3000 = 2000( 1+ i)5

( 1+i)5 = 3000 / 2000 = 1.5

1 + i = 5th root of 1.5 = 1.0844

i = 1.0844 - 1 = 0.0844 or 8.4 percent